# Island of acoustic black hole in Schwarzschild spacetime

### v1

@@ I Introduction
Recently, important progresses have been made in solving the black hole information paradox at the semiclassical level, based on the holographic calculation of entanglement entropy, i.e., the Ryu-Takayanagi (RT) formula Ryu:2006bv ; Hubeny:2007xt , and its quantum extension, i.e., the quantum extremal surface (QES) formula Faulkner:2013ana . The new approach proposed is called the island formula, which was shown can successfully reproduce the Page curve by calculating the entanglement entropy of Hawking radiation during black hole evaporation Penington:2019npb ; Almheiri:2019psf ; Almheiri:2019hni ; Almheiri:2020cfm . The formula was later confirmed by using the replica method for the gravitational path integral, where the emergence of new saddles, known as replica wormholes, lead to islands, which are regions located inside the black hole and belong to part of the entanglement wedge of the external radiation region Penington:2019kki ; Almheiri:2019qdq . The island formula explicitly states the transition in entanglement entropy, that leads to the Page curve. Explicitly, in the initial no-island phase, the entropy grows monotonically with time. After a critical point (the Page time), a new island phase dominates and makes the entropy stable. This transition ensures that the total entanglement entropy of the system follows the Page curve. The island formula can be explicitly expressed as
@@
where $R$ is the radiation region and $\partial I$ is the boundary of the island, and the area-like divergences of the entanglement entropy of matter fields that depend on the short-distance cutoff have been absorbed into the renormalized Newton’s constant $\tilde{G}_{N}$ Bombelli:1986rw ; Srednicki:1993im ; Susskind:1994sm . So far, the island formula has been successfully applied to resolve the information paradox in two dimensional gravity and has subsequently been generalized to other lower and higher dimensional black holes, see, e.g., Hollowood:2020cou ; Goto:2020wnk ; Anegawa:2020ezn ; Gautason:2020tmk ; Hartman:2020swn ; Wang:2021mqq ; Wang1:2021woy ; Li:2021lfo ; Almheiri:2019psy ; He:2021mst ; Alishahiha:2020qza ; Hashimoto:2020cas ; Krishnan:2020oun ; Hu:2022zgy ; Gan:2022jay ; Du:2022vvg ; Ling:2020laa ; Wang:2021woy ; Miao:2023unv ; Guo:2023fly ; Guo:2023gfa ; Kim:2021gzd ; Tong:2023nvi ; Du:2025kcx ; Ahn:2021chg ; HosseiniMansoori:2022hok .
@@
In the present paper, we will instead investigate the information paradox in analogue gravity. The motivations are the follows. Firstly, the typical feature of analogue gravity is the emergence of curved geometry, especially the horizon, which can produce analogue Hawking radiation composed of quantized phonons Barcelo:2005fc . Therefore, it was originally proposed to mimic curved geometries and quantum effects in curved spacetime. For example, Unruh first proposed the acoustic black hole from nonrelativistic fluids in flat spacetime to mimic the near horizon geometry and Hawking radiation of black hole Unruh:1980cg . Since then, analogue gravity have been successfully realized in various condensed matter systems such as the Bose-Einstein condensation Lahav:2009wx , the optical media Drori:2018ivu and many others (see the review Barcelo:2005fc on analogue gravity models). The advantage of analogue gravity is that they provide alternative experimental tools in lab for simulating quantum effects of in curved spacetime like the Hawking radiation of black holes Steinhauer:2015saa , which are inaccessible by current astrophysical observations due to technological limitations and precision requirements. Secondly, the analogue gravity is not merely an analogy. Using the fluid/gravity duality, it was first shown in Ge:2015uaa that the acoustic black hole formed in fluid at the radial cutoff surface can be mapped to a real black brane in asymptotically AdS spacetime at the dynamical level, the Hawking-like temperature of the acoustic black hole is connected to the Hawking temperture of the AdS black brane, and moreover, the phonon is dual to the sound channel of the quasinormal mode in the bulk AdS black brane. The dynamical connection between acoustic black hole in fluid and black hole has also been extended to the black D3 brane in flat spacetime Yu:2017bnu . Thirdly, apart from systems in flat spacetime, analogue gravity can also appear in curved spacetime, e.g., acoustic black holes have been extended in general curved spacetimes from relativistic Gross-Pitaevskii (GP) theory and Yang-Mills (YM) theory, where a real black hole is surrounded by an acoustic horizon, which is equivalent to a ‘dumb hole’ from which neither light nor sound can escape Ge:2019our . Thus the analogue gravity in curved spacetime can be used to describe real black holes surrounded by environments such as some fluid accreting on to a black hole Das:2004zm . Later works have discussed effects such as black hole shadows, quasinormal modes, and analogue Hawking radiation in this class of analogue gravity Wang:2019zqw ; Guo:2020blq ; Ling:2021vgk ; Qiao:2021trw ; Wang:2022gbl ; Parvizi:2023foz .

### v2

@@ I Introduction
Recently, important progresses have been made in solving the black hole information paradox at the semiclassical level, based on the holographic calculation of entanglement entropy, i.e., the Ryu-Takayanagi (RT) formula Ryu:2006bv ; Hubeny:2007xt , and its quantum extension, i.e., the quantum extremal surface (QES) formula Faulkner:2013ana . The new approach proposed is called the island formula, which was shown can successfully reproduce the Page curve by calculating the entanglement entropy of Hawking radiation during black hole evaporation Penington:2019npb ; Almheiri:2019psf ; Almheiri:2019hni ; Almheiri:2020cfm . The formula was later confirmed by using the replica method for the gravitational path integral, where the emergence of new saddles, known as replica wormholes, lead to islands, which are regions located inside the black hole and belong to part of the entanglement wedge of the external radiation region Penington:2019kki ; Almheiri:2019qdq ; Yu:2025tid . The island formula explicitly states the transition in entanglement entropy, that leads to the Page curve. Explicitly, in the initial no-island phase, the entropy grows monotonically with time. After a critical point (the Page time), a new island phase dominates and makes the entropy stable. This transition ensures that the total entanglement entropy of the system follows the Page curve. The island formula can be explicitly expressed as
@@
where $R$ is the radiation region and $\partial I$ is the boundary of the island, and the area-like divergences of the entanglement entropy of matter fields that depend on the short-distance cutoff have been absorbed into the renormalized Newton’s constant $\tilde{G}_{N}$ Bombelli:1986rw ; Srednicki:1993im ; Susskind:1994sm . So far, the island formula has been successfully applied to resolve the information paradox in two dimensional gravity and has subsequently been generalized to other lower and higher dimensional black holes, see, e.g., Hollowood:2020cou ; Goto:2020wnk ; Anegawa:2020ezn ; Gautason:2020tmk ; Hartman:2020swn ; Wang:2021mqq ; Wang1:2021woy ; Li:2021lfo ; Almheiri:2019psy ; He:2021mst ; Alishahiha:2020qza ; Hashimoto:2020cas ; Krishnan:2020oun ; Hu:2022zgy ; Gan:2022jay ; Du:2022vvg ; Ling:2020laa ; Wang:2021woy ; Miao:2023unv ; Guo:2023fly ; Guo:2023gfa ; Kim:2021gzd ; Tong:2023nvi ; Du:2025kcx ; Ahn:2021chg ; HosseiniMansoori:2022hok ; RoyChowdhury:2022awr ; Yu:2024fks ; Espindola:2022fqb .
@@
In the present paper, we will instead investigate the information paradox in analogue gravity. The motivations are the follows. Firstly, the typical feature of analogue gravity is the emergence of curved geometry, especially the horizon, which can produce analogue Hawking radiation composed of quantized phonons Barcelo:2005fc . Therefore, it was originally proposed to mimic curved geometries and quantum effects in curved spacetime. For example, Unruh first proposed the acoustic black hole from nonrelativistic fluids in flat spacetime to mimic the near horizon geometry and Hawking radiation of black hole Unruh:1980cg . Since then, analogue gravity have been successfully realized in various condensed matter systems such as the Bose-Einstein condensation Lahav:2009wx , the optical media Drori:2018ivu and many others (see the review Barcelo:2005fc on analogue gravity models). The advantage of analogue gravity is that they provide alternative experimental tools in lab for simulating quantum effects of in curved spacetime like the Hawking radiation of black holes Steinhauer:2015saa , which are inaccessible by current astrophysical observations due to technological limitations and precision requirements. [^1] There are also relevant studies discussing the simulation of physical phenomena such as cosmic inflation and particle pair production through analogue gravity Fischer:2004bf ; Fedichev:2003id ; Cha:2016esj . Secondly, the analogue gravity is not merely an analogy. Using the fluid/gravity duality, it was first sown in Ge:2015uaa that the acoustic black hole formed in fluid at the radial cutoff surface can be mapped to a real black brane in asymptotically AdS spacetime at the dynamical level, the Hawking-like temperature of the acoustic black hole is connected to the Hawking temperture of the AdS black brane, and moreover, the phonon is dual to the sound channel of the quasinormal mode in the bulk AdS black brane. The dynamical connection between acoustic black hole in fluid and black hole has also been extended to the black D3 brane in flat spacetime Yu:2017bnu . Thirdly, apart from systems in flat spacetime, analogue gravity can also appear in curved spacetime, e.g., acoustic black holes have been extended in general curved spacetimes from relativistic Gross-Pitaevskii (GP) theory and Yang-Mills (YM) theory, where a real black hole is surrounded by an acoustic horizon, which is equivalent to a ‘dumb hole’ from which neither light nor sound can escape Ge:2019our . Thus the analogue gravity in curved spacetime can be used to describe real black holes surrounded by environments such as some fluid accreting on to a black hole Das:2004zm . Later works have discussed effects such as black hole shadows, quasinormal modes, and analogue Hawking radiation in this class of analogue gravity Wang:2019zqw ; Guo:2020blq ; Ling:2021vgk ; Qiao:2021trw ; Wang:2022gbl ; Parvizi:2023foz .